OpenStudy (anonymous):

Adding and Subtracting with unlike denominators Add or Subtract. Write answer in simplest form

OpenStudy (anonymous):

I'm posting it one moment

OpenStudy (anonymous):

\[\frac{ 3 }{ 5 } + \frac{ 2 }{ 7 }\]

OpenStudy (anonymous):

you need a common denominator. so multiply both numbers by 1. \[(7/7)*(3/5)+(5/5)*(2/7)\]

OpenStudy (anonymous):

Once the denominator is the same just add the numbers on top. \[= (21/35) + (10/35) = 31/35 \]

OpenStudy (anonymous):

Does that make sense?

OpenStudy (anonymous):

\[ \frac{ 31 }{ 35 }, \frac{ 1 }{ 7 }, \frac{ 5 }{ 7 }. \frac{ 6 }{ 35 } \]

OpenStudy (anonymous):

Yes it makes lots of sense thank you

OpenStudy (anonymous):

So you're correct

OpenStudy (anonymous):

any other questions?